# Magnitude homology of real hyperplane arrangements

### 04 Apr 2026 v1

@@ 3.3. Examples
We use the notion $\mathcal{A}(n,k)$ from Grünbaum’s catalogue [14].
@@
The near-pencil $\mathcal{A}(n,0)$ has tope graph $C_{2(n-1)}\Box K_{2}$ and hence
@@
Note that $\mathcal{A}(6,1)$ is the Coxeter arrangement of type $A_{3}$ or $D_{3}$, and $\mathcal{A}(9,1)$ is the Coxeter arrangement of type $B_{3}$, which can be computed from Proposition 3.5. In what follows, $\Phi_{k}(q)\in\mathbb{Z}[q]$ is the $k$-th cyclotomic polynomial.
@@
More examples are in the appendix.
4. Magnitude homology of tope graphs
@@
We continue denoting invariants of $\mathcal{T}(\mathcal{A})$ simply by invariants of $\mathcal{A}$, such as $MH_{k,\ell}(\mathcal{A}):=MH_{k,\ell}(\mathcal{T}(\mathcal{A}))$, $\beta_{k,\ell}(\mathcal{A}):=\mathrm{rank}MH_{k,\ell}(\mathcal{A})$ and $\chi_{\ell}(\mathcal{A})=\sum_{k}(-1)^{k}\beta_{k,\ell}(\mathcal{A})$. Also recall the notation $c^{X}=\#\mathrm{Ch}(\mathcal{A}^{X})$ for $X\in L(\mathcal{A})$.
4.1. Small length
Let us start our study of $MH_{k,\ell}(\mathcal{A})$ from investigating small length $\ell\leq 2$.
@@ 4.1. Small length
Let $\mathcal{A}$ be a real central hyperplane arrangement. Then the only nontrivial $MH_{k,\ell}(\mathcal{A})$ for $\ell\leq 2$ are as follows.
@@
(1) and (2) follow from (1) and (2) of Proposition 2.8 and noting that an edge in $\mathcal{T}(\mathcal{A})$ corresponds to a chamber of restriction $\mathcal{A}^{H}$ for some $H\in\mathcal{A}$.
For (3), since $m_{\mathcal{T}(\mathcal{A})}\geq 3$, we may apply Theorem 2.20 for $\ell=2<m_{\mathcal{T}(\mathcal{A})}$.
$MH_{2,2}(\mathcal{A})\cong\bigoplus_{\vec{a}\in P_{1,2}(\mathcal{A})\cup P_{2,2}(\mathcal{A})}MH_{2,\vec{a}}(\mathcal{A}).$
By Theorem 2.23, the framed magnitude homology of $\mathcal{A}$ can be computed from interval posets. For $\vec{a}=(a_{0},a_{1})\in P_{1,2}(\mathcal{A})$, where $a_{0},a_{1}$ are chambers at distance $2$, we may suppose $S(a_{0},a_{1})=\{H_{1},H_{2}\}$ and let $X=H_{1}\cap H_{2}\in L_{2}(\mathcal{A})$. If $\mathcal{A}_{X}=2$, then the interval poset $(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}$ consists of two incomparable chambers between $a_{0}$ and $a_{1}$. If $\mathcal{A}_{X}>2$, then $(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}$ is a singleton. Since $MH_{2,(a_{0},a_{1})}(\mathcal{A})\cong\widetilde{H}_{0}(C(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})})$, taking reduced homology, only the first type $(a_{0},a_{1})$ with $\mathcal{A}_{X}=2$ contributes to the magnitude homology a copy of $\mathbb{Z}$. Then the total contribution to $MH_{2,2}(\mathcal{A})$ from $P_{1,2}(\mathcal{A})$ is $4\sum_{X\in L_{2}(\mathcal{A}),\#\mathcal{A}_{X}=2}c^{X}$ copies of $\mathbb{Z}$.
Similarly, for $\vec{a}=(a_{0},a_{1},a_{2})\in P_{2,2}(\mathcal{A})$, if $a_{0}\neq a_{2}$, then $\vec{a}$ is geodesic and hence $\varphi(\vec{a})\neq\vec{a}$, making the $\vec{a}$-framed magnitude homology trivial. So we only need to consider $\vec{a}=(a_{0},a_{1},a_{0})\in P_{2,2}(\mathcal{A})$, where $a_{0}$ and $a_{1}$ are adjacent. The interval posets $(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}$ and $(a_{1},a_{0})_{\mathcal{T}(\mathcal{A})}$ are empty. Then $MH_{2,(a_{0},a_{1},a_{0})}(\mathcal{A})\cong\widetilde{H}_{-2}(C(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}\otimes C(a_{1},a_{0})_{\mathcal{T}(\mathcal{A})})\cong\mathbb{Z}$. Each such $(a_{0},a_{1},a_{0})$ corresponds to a directed edge of $\mathcal{T}(\mathcal{A})$. Then the total contribution to $MH_{2,2}(\mathcal{A})$ from $P_{2,2}(\mathcal{A})$ is $\beta_{1,1}(\mathcal{A})$ copies of $\mathbb{Z}$. Now (3) is proved.
To finish the proof, we need to show other $MH_{k,\ell}(\mathcal{A})$ for $\ell\leq 2$ are trivial. By Proposition 2.8 (3), we only need to prove $MH_{1,2}(\mathcal{A})=0$. We use Theorem 2.20 again
$MH_{1,2}(A)\cong\bigoplus_{\vec{a}\in P_{1,2}(\mathcal{A})}MH_{1,\vec{a}}(\mathcal{A}).$
Consider $\vec{a}=(a_{0},a_{1})\in P_{1,2}(\mathcal{A})$, the interval poset $(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}$ cannot be empty. Hence $MH_{1,(a_{0},a_{1})}(\mathcal{A})\cong\widetilde{H}_{-1}(C(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})})=0$. ∎

### 01 May 2026 v3

@@ 3.3. Examples
We use the notation $\mathcal{A}(n,k)$ from Grünbaum’s catalogue [15].
@@
The near-pencil $\mathcal{A}(n,0)$ has the tope graph $C_{2(n-1)}\Box K_{2}$ and hence
@@
Note that $\mathcal{A}(6,1)$ is the Coxeter arrangement of type $A_{3}$ or $D_{3}$, and $\mathcal{A}(9,1)$ is the Coxeter arrangement of type $B_{3}$. Therefore, the magnitudes of both arrangements can be computed from Proposition 3.5. In what follows, $\Phi_{k}(q)\in\mathbb{Z}[q]$ is the $k$-th cyclotomic polynomial.
@@
Additional examples are given in the appendix.
4. Computations for rank 3 arrangements
4.1. Magnitude of rank 3 arrangements
Using the face decomposition formula (Theorem 3.7), we can explicitly compute the magnitude of rank $3$ arrangements.
Let $\mathcal{A}$ be an arrangement of rank $3$ with $n=\#\mathcal{A}$. For $k\geq 2$, we write $n_{k}$ for the number of faces $F\in\mathcal{F}(\mathcal{A})$ with $\operatorname{codim}F=2$ and $\mathcal{A}_{F}=k$.
Theorem 4.1.
Let $\mathcal{A}$ be an arrangement of rank $3$ with $\mathcal{A}=n$. Then we have
$\mathrm{Mag}(\mathcal{A})=\dfrac{1}{1+q^{n}}\biggl(\#\mathrm{Ch}(\mathcal{A})-\sum_{k\geq 2}\dfrac{2kn_{k}q(1-q^{k-1})}{(1+q)(1-q^{k})}\biggr).$
Proof.
Let $\mathcal{F}_{d}(\mathcal{A})$ denote the set of faces of codimension $d$ of $\mathcal{A}$. By Theorem 3.7, we have
$\mathrm{Mag}(\mathcal{A})$ ${}=\sum_{F\in\mathcal{F}_{0}(\mathcal{A})}1-\sum_{F\in\mathcal{F}_{1}(\mathcal{A})}\dfrac{2q}{1+q}+\sum_{F\in\mathcal{F}_{2}(\mathcal{A})}\dfrac{2\#\mathcal{A}_{F}q^{\#\mathcal{A}_{F}}(1-q)}{(1+q)(1-q^{\#\mathcal{A}_{F}})}-q^{n}\mathrm{Mag}(\mathcal{A},q)$
${}=\#\mathrm{Ch}(\mathcal{A})-\#\mathcal{F}_{1}(\mathcal{A})\cdot\dfrac{2q}{1+q}+\sum_{k\geq 2}\dfrac{2kn_{k}q^{k}(1-q)}{(1+q)(1-q^{k})}-q^{n}\mathrm{Mag}(\mathcal{A},q).$
For each $F\in\mathcal{F}_{2}(\mathcal{A})$, the number of faces $G\in\mathcal{F}_{1}(\mathcal{A})$ of codimension $1$ adjacent to $F$ is exactly $2\#\mathcal{A}_{F}$. Conversely, for each $G\in\mathcal{F}_{1}(\mathcal{A})$, there are exactly two faces $F\in\mathcal{F}_{2}(\mathcal{A})$ of codimension $2$ adjacent to $G$. Therefore we have
$\mathcal{F}_{1}(\mathcal{A})=\dfrac{1}{2}\sum_{F\in\mathcal{F}_{2}(\mathcal{A})}2\#\mathcal{A}_{F}=\sum_{k\geq 2}kn_{k}.$
Using this identity, we obtain
$(1+q^{n})\mathrm{Mag}(\mathcal{A})$ ${}=\#\mathrm{Ch}(\mathcal{A})-\sum_{k\geq 2}\dfrac{2kn_{k}q}{1+q}+\sum_{k\geq 2}\dfrac{2kn_{k}q^{k}(1-q)}{(1+q)(1-q^{k})}$
${}=\#\mathrm{Ch}(\mathcal{A})-\sum_{k\geq 2}\dfrac{2kn_{k}q(1-q^{k-1})}{(1+q)(1-q^{k})}.$
This completes the proof. ∎
Example 4.2.
The arrangement $\mathcal{A}(6,1)$ from Grünbaum’s catalogue [15] satisfies
$\mathrm{Ch}(\mathcal{A}(6,1))=24\quad\text{and}\quad n_{k}=\begin{cases}6,&\text{ if }k=2;\\ 8,&\text{ if }k=3;\\ 0,&\text{ if }k\geq 4.\end{cases}$
By Theorem 4.1, its magnitude is given by
$\mathrm{Mag}(\mathcal{A}(6,1))$ ${}=\dfrac{1}{1+q^{6}}\biggl(24-\dfrac{24q}{(1+q)^{2}}-\dfrac{48q}{1+q+q^{2}}\biggr)$
${}=\frac{24}{(1+q)^{2}(1+q+q^{2})(1+q^{2})}.$
This coincides with the formula given in Example 3.13.
4.2. Alternating sign property
In all the examples shown in Appendix A, the coefficients in the series expansion of the magnitude of each arrangement seem to have alternating signs. That is, if we write
$\mathrm{Mag}(\mathcal{A})=\sum_{\ell=0}^{\infty}c_{\ell}q^{\ell},$
then $(-1)^{\ell}c_{\ell}\geq 0$ appears to hold. Here we prove that if $\operatorname{rank}\mathcal{A}\leq 3$, then $(-1)^{\ell}c_{\ell}>0$ holds for all sufficiently large $\ell$. On the other hand, in the next subsection, we construct an example for which the inequality $(-1)^{\ell}c_{\ell}\geq 0$ fails for small $\ell$.
Definition 4.3.
For each $k\geq 2$, we define a rational function $G_{k}(q)$ and a sequence $(g_{k,\ell})_{\ell\geq 0}$ by
$G_{k}(q)=\dfrac{q(1-(-q)^{k-1})}{(1-q)(1-(-q)^{k})}=\sum_{\ell=0}^{\infty}g_{k,\ell}q^{\ell}.$
Theorem 4.1 implies that, for an arrangement $\mathcal{A}$ of rank $3$ with $\mathcal{A}=n$, we have
(4.1) $\mathrm{Mag}(\mathcal{A},-q)=\dfrac{1}{1+(-q)^{n}}\biggl(\#\mathrm{Ch}(\mathcal{A})+\sum_{k\geq 2}2kn_{k}G_{k}(q)\biggr).$
Lemma 4.4.
Let $\mathcal{A}$ be an arrangement of rank $3$. Let $n_{k}$ be the number of faces $F\in\mathcal{F}(\mathcal{A})$ with $\operatorname{codim}F=2$ and $\mathcal{A}_{F}=k$. Then we have $n_{2}>0$.
Proof.
We may assume that $\mathcal{A}$ is essential. Consider the zonotope $Z(\mathcal{A})$ of $\mathcal{A}$, which is a polytope in $\mathbb{R}^{3}$. The facets of $Z(\mathcal{A})$ are polygons with an even number of vertices. The number of facets with $2k$ vertices is exactly $n_{k}$. If $n_{2}=0$, then every facet of $Z(\mathcal{A})$ must have at least $6$ vertices. However, Euler’s polyhedral formula shows that this is impossible. Therefore we have $n_{2}>0$. ∎
Theorem 4.5.
Let $\mathcal{A}$ be an arrangement with $\operatorname{rank}\mathcal{A}\leq 3$. Let $c_{\ell}$ denote the coefficient of $q^{\ell}$ in the series expansion of $\mathrm{Mag}(\mathcal{A})$:
$\mathrm{Mag}(\mathcal{A})=\sum_{\ell=0}^{\infty}c_{\ell}q^{\ell}.$
Then we have $(-1)^{\ell}c_{\ell}>0$ for all sufficiently large $\ell$.
Proof.
When $\operatorname{rank}\mathcal{A}\leq 2$, the claim can be checked directly. Therefore, we may assume that $\operatorname{rank}\mathcal{A}=3$. Let $\mathcal{A}=n$. Write $n_{k}$ for the number of faces $F\in\mathcal{F}(\mathcal{A})$ with $\operatorname{codim}F=2$ and $\mathcal{A}_{F}=k$. Define $(e_{\ell})_{\ell\geq 0}$ by
$\mathrm{Mag}(\mathcal{A},-q)=\dfrac{1}{1+(-q)^{n}}\biggl(\#\mathrm{Ch}(\mathcal{A})+\sum_{k\geq 2}2kn_{k}G_{k}(q)\biggr)=\sum_{\ell=0}^{\infty}e_{\ell}q^{\ell}.$
It suffices to show that $e_{\ell}>0$ for all sufficiently large $\ell$. Put
$H(q):=\#\mathrm{Ch}(\mathcal{A})+\sum_{k=2}^{\infty}2kn_{k}G_{k}(q),\qquad H(q)=\sum_{\ell\geq 0}a_{\ell}q^{\ell}.$
We first record two elementary facts about the coefficients $g_{k,\ell}$ of $G_{k}$. If $k$ is even, then
$G_{k}(q)=\frac{q(1+q^{k-1})}{(1-q)(1-q^{k})}=\frac{q}{1-q}\sum_{j\geq 0}\left(q^{kj}+q^{kj+k-1}\right).$
Hence $g_{k,\ell}$ is non-negative and non-decreasing in $\ell$. In particular, for $k=2$ we have $g_{2,\ell}=\ell$. If $k$ is odd, then
$G_{k}(q)=\frac{q(1-q^{k-1})}{(1-q)(1+q^{k})}=\frac{q(1+q+\cdots+q^{k-2})}{1+q^{k}}.$
Thus the coefficient sequence $(g_{k,\ell})_{\ell\geq 0}$ is periodic and bounded. Since $n_{2}>0$ by Lemma 4.4, there is a constant $C>0$ such that $a_{\ell}\geq 4n_{2}\ell-C$ for all $\ell\geq 0$.
We now distinguish the parity of $n$. First suppose that $n$ is odd. Then $1+(-q)^{n}=1-q^{n}$, so
$\mathrm{Mag}(\mathcal{A},-q)=\frac{H(q)}{1-q^{n}}=H(q)\sum_{j\geq 0}q^{nj}.$
Therefore, with $N=\lfloor\ell/n\rfloor$,
$e_{\ell}=\sum_{j=0}^{N}a_{\ell-nj}\geq\sum_{j=0}^{N}\bigl(4n_{2}(\ell-nj)-C\bigr).$
The right-hand side is a polynomial of degree $1$ in $\ell$ with leading coefficient $4n_{2}(N+1)>0$, so it tends to $\infty$ as $\ell\to\infty$. Hence $e_{\ell}>0$ for all sufficiently large $\ell$ in this case.
Next suppose that $n$ is even. Then $1+(-q)^{n}=1+q^{n}$ and
$\frac{1}{1+q^{n}}=\sum_{j\geq 0}(-1)^{j}q^{nj}.$
For even $k$, let $b_{k,\ell}$ be the coefficient of $q^{\ell}$ in the series expansion of $\frac{G_{k}(q)}{1+q^{n}}$. Then we have
$b_{k,\ell}=\sum_{j=0}^{\lfloor\ell/n\rfloor}(-1)^{j}g_{k,\ell-nj}.$
Since $(g_{k,\ell})_{\ell}$ is non-negative and non-decreasing, pairing adjacent terms gives $b_{k,\ell}\geq 0$ for every $\ell$. For $k=2$, using $g_{2,m}=m$, we also get $b_{2,\ell}\to+\infty$. We claim that the odd $k$ terms, after division by $1+q^{n}$, are bounded. Indeed, for odd $k$ and even $n$,
$\frac{G_{k}(q)}{1+q^{n}}=\frac{q(1+q+\cdots+q^{k-2})}{(1+q^{k})(1+q^{n})}.$
Moreover $\gcd(1+q^{k},1+q^{n})=1$, because $k$ is odd and $n$ is even. Hence the denominator has only simple roots of unity, and the coefficient sequence is bounded. The coefficient sequence of $\mathrm{Ch}(\mathcal{A})/(1+q^{n})$ is also bounded. Thus there is a constant $D>0$ such that the total contribution of $\mathrm{Ch}(\mathcal{A})$ and all odd $k$ terms to $e_{\ell}$ is at least D-D for every $\ell$. Combining these observations, and using Lemma 4.4, we obtain
$e_{\ell}\geq 4n_{2}b_{2,\ell}-D.$
Since $b_{2,\ell}\to+\infty$, the right-hand side is positive for all sufficiently large $\ell$. This proves the theorem. ∎
4.3. Arrangements with non-alternating magnitude
In this section, we use the construction of Burr–Grünbaum–Sloane [5] for the orchard problem to construct an example of an arrangement whose magnitude does not have the alternating-sign property.
Recall the formula (4.1) for $\mathrm{Mag}(\mathcal{A},-q)$. The series expansion of $G_{3}(q)$ is given by
$G_{3}(q)=\dfrac{q}{1-q+q^{2}}=q+q^{2}-q^{4}-q^{5}+q^{7}+q^{8}-\cdots.$
This suggests that for an arrangement with large $n_{3}$, the series expansion of $\mathrm{Mag}(\mathcal{A},-q)$ might have negative coefficients. This motivates the construction given below.
Definition 4.6.
Let $S\subset\mathbb{R}^{2}$ be a finite set of points in $\mathbb{R}^{2}$. For each $k\geq 2$, define $m_{k}$ to be the number of lines that contain exactly $k$ points from $S$. Define an arrangement $\mathcal{A}(S)$ in $\mathbb{R}^{3}$ by
$\mathcal{A}(S)=\{(x,y,1)^{\perp}\mid(x,y)\in S\}.$
The arrangement $\mathcal{A}(S)$ is essential unless $S$ is contained in a single line. It satisfies $n_{k}=2m_{k}$ for $k\geq 2$.
We use the following result to construct an $S$ with large $m_{3}$.
Theorem 4.7 ([5, Theorem 1], see also [14, Proposition 2.6]).
For any $N\geq 3$, there exists a finite set $S\subset\mathbb{R}^{2}$ with #S=N\#S=N such that
$m_{k}=\begin{cases}N-1-2\cdot\mathbf{1}_{3\mid N},&\text{ if }k=2;\\ \biggl\lfloor\dfrac{N(N-3)}{6}+1\biggr\rfloor,&\text{ if }k=3;\\ 0,&\text{ if }k\geq 4,\end{cases}$
where $\mathbf{1}_{3\mid N}$ is equal to $1$ when $3$ divides $N$ and 0 otherwise.
Example 4.8.
As an example of Theorem 4.7, we construct a set of points $S\subset\mathbb{R}^{2}$ with #S=18\#S=18 such that
$m_{k}=\begin{cases}15,&\text{ if }k=2;\\ 46,&\text{ if }k=3;\\ 0,&\text{ if }k\geq 4.\end{cases}$
Let $X$ be the set of non-singular points of the real projective algebraic curve
$Y^{2}=X^{2}(X-Z).$
Then the points of $X$ naturally form a group isomorphic to $\mathbb{R}/\mathbb{Z}$. An explicit parametrization is given by
$\phi(t)=\bigl[\cot^{2}(\pi t)+1:\cot(\pi t)(\cot^{2}(\pi t)+1):1\bigr]\quad(t\in\mathbb{R}/\mathbb{Z}).$
Let $P_{j}\in\mathbb{R}^{2}$ be the image of $\phi(j/18)$ under the projective transformation $X:Y:Z]\mapsto[X:Y:X+Y+Z]$. Then $P_{i},P_{j},P_{k}$ lie on a single line if and only if $i+j+k\equiv 0\pmod{18}$. Define $S\subset\mathbb{R}^{2}$ by
$S=\{P_{j}\mid j=0,1,\dots,17\}.$
There are exactly $46$ triples $\{i,j,k\}\subset\mathbb{Z}/18\mathbb{Z}$ such that $i+j+k\equiv 0\pmod{18}$, so we have $m_{3}=46$. Also, $m_{k}=0$ for $k\geq 4$. Since $m_{2}+3m_{3}=\binom{18}{2}$ by double-counting, we have $m_{2}=15$.
Example 4.9.
We give an explicit construction of an arrangement $\mathcal{A}$ such that the series expansion of $\mathrm{Mag}(\mathcal{A})$ does not have alternating signs. Let $S\subset\mathbb{R}^{2}$ be a finite set as in Example 4.8. Then the associated arrangement $\mathcal{A}(S)$ satisfies
$\mathcal{A}(S)=18,\quad n_{k}=\begin{cases}30,&\text{ if }k=2;\\ 92,&\text{ if }k=3;\\ 0,&\text{ if }k\geq 4,\end{cases}\quad\text{and}\quad\#\mathrm{Ch}(\mathcal{A})=216.$
Figure 2 shows the projectivization of $\mathcal{A}(S)$. By Theorem 4.1, the magnitude of $\mathcal{A}(S)$ is given by
$\mathrm{Mag}(\mathcal{A})$ ${}=\dfrac{1}{1+q^{18}}\biggl(216-\dfrac{120q}{(1+q)^{2}}-\dfrac{552q}{1+q+q^{2}}\biggr)$
${}=216-672q+792q^{2}-360q^{3}-72q^{4}-48q^{5}+720q^{6}-1392q^{7}+1512q^{8}-\cdots.$
In particular, the coefficient of $q^{4}$ is 72<0-72<0.
Figure 2. An arrangement whose magnitude does not have alternating signs
5. Magnitude homology of arrangements
@@
We continue to denote invariants of $\mathcal{T}(\mathcal{A})$ simply by the corresponding invariants of $\mathcal{A}$, such as $MH_{k,\ell}(\mathcal{A}):=MH_{k,\ell}(\mathcal{T}(\mathcal{A}))$, $\beta_{k,\ell}(\mathcal{A}):=\operatorname{rank}MH_{k,\ell}(\mathcal{A})$, and $\chi_{\ell}(\mathcal{A})=\sum_{k}(-1)^{k}\beta_{k,\ell}(\mathcal{A})$. Also recall the notation $c^{X}=\#\mathrm{Ch}(\mathcal{A}^{X})$ for $X\in L(\mathcal{A})$.
5.1. Small lengths
We begin our study of $MH_{k,\ell}(\mathcal{A})$ by investigating small lengths $\ell\leq 2$.
@@ 5.1. Small lengths
Let $\mathcal{A}$ be an arrangement. Then the only nontrivial $MH_{k,\ell}(\mathcal{A})$ for $\ell\leq 2$ are as follows.
@@
(1) and (2) follow from (1) and (2) of Proposition 2.8, together with the observation that an edge in $\mathcal{T}(\mathcal{A})$ corresponds to a chamber of the restriction $\mathcal{A}^{H}$ for some $H\in\mathcal{A}$.
For (3), since $m_{\mathcal{T}(\mathcal{A})}\geq 3$, we may apply Theorem 2.22 for $\ell=2<m_{\mathcal{T}(\mathcal{A})}$:
$MH_{2,2}(\mathcal{A})\cong\bigoplus_{\vec{a}\in P_{1,2}(\mathcal{A})\cup P_{2,2}(\mathcal{A})}MH^{\mathrm{gs}}_{2,\vec{a}}(\mathcal{A}).$
By Theorem 2.25, the framed magnitude homology of $\mathcal{A}$ can be computed from interval posets. For $\vec{a}=(a_{0},a_{1})\in P_{1,2}(\mathcal{A})$, where $a_{0},a_{1}$ are chambers at distance $2$, we may suppose $S(a_{0},a_{1})=\{H_{1},H_{2}\}$ and let $X=H_{1}\cap H_{2}\in L_{2}(\mathcal{A})$. If $\mathcal{A}_{X}=2$, then the interval poset $(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}$ consists of two incomparable chambers between $a_{0}$ and $a_{1}$. If $\mathcal{A}_{X}>2$, then $(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}$ is a singleton. Since $MH^{\mathrm{gs}}_{2,(a_{0},a_{1})}(\mathcal{A})\cong\widetilde{H}_{0}(C(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})})$, taking reduced homology shows that only the first type of $(a_{0},a_{1})$, with $\mathcal{A}_{X}=2$, contributes a copy of $\mathbb{Z}$ to the magnitude homology. Hence the total contribution to $MH_{2,2}(\mathcal{A})$ from $P_{1,2}(\mathcal{A})$ is $4\sum_{X\in L_{2}(\mathcal{A}),\#\mathcal{A}_{X}=2}c^{X}$ copies of $\mathbb{Z}$.
Similarly, for $\vec{a}=(a_{0},a_{1},a_{2})\in P_{2,2}(\mathcal{A})$, if $a_{0}\neq a_{2}$, then $\vec{a}$ is geodesic and hence $\varphi(\vec{a})\neq\vec{a}$, making the $\vec{a}$-framed magnitude homology trivial. Thus we only need to consider $\vec{a}=(a_{0},a_{1},a_{0})\in P_{2,2}(\mathcal{A})$, where $a_{0}$ and $a_{1}$ are adjacent. The interval posets $(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}$ and $(a_{1},a_{0})_{\mathcal{T}(\mathcal{A})}$ are empty. Then $MH^{\mathrm{gs}}_{2,(a_{0},a_{1},a_{0})}(\mathcal{A})\cong\widetilde{H}_{-2}(C(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}\otimes C(a_{1},a_{0})_{\mathcal{T}(\mathcal{A})})\cong\mathbb{Z}$. Each such $(a_{0},a_{1},a_{0})$ corresponds to a directed edge of $\mathcal{T}(\mathcal{A})$. Hence the total contribution to $MH_{2,2}(\mathcal{A})$ from $P_{2,2}(\mathcal{A})$ is $\beta_{1,1}(\mathcal{A})$ copies of $\mathbb{Z}$. This proves (3).
To finish the proof, we need to show that all other $MH_{k,\ell}(\mathcal{A})$ for $\ell\leq 2$ are trivial. By Proposition 2.8 (3), we only need to prove $MH_{1,2}(\mathcal{A})=0$. We use Theorem 2.22 again:
$MH_{1,2}(\mathcal{A})\cong\bigoplus_{\vec{a}\in P_{1,2}(\mathcal{A})}MH^{\mathrm{gs}}_{1,\vec{a}}(\mathcal{A}).$
Consider $\vec{a}=(a_{0},a_{1})\in P_{1,2}(\mathcal{A})$. The interval poset $(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}$ cannot be empty. Hence $MH^{\mathrm{gs}}_{1,(a_{0},a_{1})}(\mathcal{A})\cong\widetilde{H}_{-1}(C(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})})=0$. ∎